### v1

# Diversity Enhances an LLM’s Performance in RAG and Long-context Task

Abstract
The rapid advancements in large language models (LLMs) have highlighted the challenge of context window limitations, primarily due to the quadratic time complexity of the self-attention mechanism ($O(N^{2})$, where $N$ denotes the context window length). This constraint impacts tasks such as retrieval-augmented generation (RAG) in question answering (Q&A) and long context summarization. A common approach involves selecting content with the highest similarity to the query; however, this often leads to redundancy and the exclusion of diverse yet relevant information. Building on principles from Maximal Marginal Relevance (MMR) and Farthest Point Sampling (FPS), we integrate diversity into the content selection process. Our findings reveal that incorporating diversity substantially increases the recall of selecting relevant sentences or chunks before LLM-based Q&A and summarization. These results highlight the importance of maintaining diversity in future LLM applications to further improve summarization and Q&A outcomes.
footnotetext: *: corresponding author
     github code: https://github.com/ZhichaoWang970201/DIVERSITY-ENHANCES-LLM
Figure 1: For both Q&A and summarization tasks, the initial dataset is divided into sentences or chunks, and corresponding embeddings are extracted. In a traditional pipeline, query embeddings are generated and used to select relevant materials to LLMs for downstream tasks. In contrast, methods like MMR and FPS incorporate diversity in a greedy manner when selecting relevant sentences. This approach increases the likelihood of including the correct answer within the chosen sentences or chunks.
1 Introduction
The remarkable success of Transformer models [1], BERT [2], and GPT [3] can be largely attributed to their robust self-attention mechanisms. However, the self-attention module’s quadratic time complexity, $O(N^{2})$, where $N$ represents the context window length, has imposed limitations on the size of the context window.
Recent advances in LLMs have partially addressed this constraint. For instance, GPT-3.5 demonstrates the capability to process context windows of up to 16,385 tokens, while GPT-4 extends this capacity to an impressive 128,000 tokens. Despite these notable improvements, the challenge of processing even longer sequences remains a critical area of research for several compelling reasons. First, many real-world applications, such as question-answering systems operating on extensive datasets, cannot accommodate entire document collections within the LLM’s context window. This limitation has led to the development of Retrieval-Augmented Generation (RAG) systems [4], which selectively retrieve and process relevant text segments for specific queries. Second, while current context window sizes may suffice for conventional Natural Language Processing (NLP) tasks, they prove inadequate for high-frequency signal processing applications. For example, audio processing and medical vibrational signal analysis often require handling data streams with sampling rates reaching one million samples per second, far exceeding current context window capabilities [5]. Furthermore, empirical studies have revealed a concerning trend: LLM performance tends to degrade as input lengths approach the maximum context window capacity, highlighting the need for more robust solutions to long-sequence processing [6].
Various strategies have been devised to address the limited context window issue in LLMs. Longformer [7] applies attention to immediate local neighbors, reducing the time complexity from $O(N^{2})$ to $O(NM)$, with $M$ representing the considered neighbors. This approach, however, necessitates significant alterations to the attention mechanism, which is not commonly adopted in contemporary LLMs such as GPT[3], Llama[8], and Gemini[9]. An alternative strategy is to expand the context window at inference [10]. Although this can mitigate the modification during the training process, it still demands changes to the attention architecture during the inference time, which is not accessible for close-source models like GPT.
Several strategies have been proposed to address the limited context window in LLM from the training perspective. The Longformer model [7] employs attention mechanisms focused on immediate local neighbors, reducing the time complexity from $O(N^{2})$ to $O(NM)$, where $M$ denotes the number of neighbors considered. However, this method requires substantial modifications to the attention mechanism, which are not widely adopted by contemporary LLMs such as GPT [3], LLaMA [8], and Gemini [9]. Another approach involves extending the context window during inference [10]. While this mitigates the need for training-time modifications, it necessitates changes to the attention mechanism at inference and full access to the model architecture—an obstacle for closed-source models like GPT.
Previous methods primarily focus on modifying LLMs to increase their context window. However, a more straightforward approach is to first select the most relevant documents while ensuring they fit within the LLM’s context window. For a given query, multiple documents are split into smaller chunks or sentences. The embeddings for both the query and the split documents are then computed. Similarity metrics, such as cosine similarity or Euclidean distance, are subsequently used to identify the most relevant sentences.
However, relying solely on the similarity between a query and segmented documents can result in overlooking critical information due to excessive focus on similar content. Previous studies have introduced greedy algorithms, such as MMR [11] and FPS [12], to improve diversity during the selection process. Related work introduced Hypothetical Document Embedding (HyDE) and LLM reranking to enhance diversity in Q&A tasks, claiming their method outperforms MMR [13, 14]. However, these studies did not address the recall of relevant documents prior to LLM generation, which is more pertinent to diversity considerations. Additionally, they did not explore various hyperparameters within MMR. Then, they have not explored the impact of reordering of selected sentences or chunks on the downstream tasks. In this paper, we aim to address this gap by conducting experiments to demonstrate the significance of diversity in long context summarization and RAG-based Q&A tasks at multiple levels: sentence-level for single documents, chunk-level across entire datasets, and sentence-level in summarization.
The contributions of this paper are summarized as follows:
  1. 1.
    We demonstrate the benefits of diversity using MMR and FPS with proper hyperparameters, i.e., $\alpha$ and $w$ on downstream tasks, including Q&A and summarization.
  2. 2.
    We discover that MMR achieves slightly better recall than FPS while maintaining significantly lower latency.
  3. 3.
    We prove the ordering selected sentences within the original document and ordering selected chunks based on the scores has the best downstream performances.
2 Methodology
In this section, we will start with a brief introduction of MMR and FPS to consider diversity during the search process. Then, the integration with LLM will be discussed.
2.1 MMR and FPS for Diversity
MMR
The concept of MMR involves selecting a subset $S$ from a large dataset $T$ [11]. MMR uses a greedy algorithm that starts with the selected set $S$ being empty and the remaining set $R$ being the entire dataset $T$. In each iteration, an element is chosen based on a locally optimal selection process, as defined in Eq. 1. The parameter $\alpha$ balances the trade-off between rewards and diversity. Let $r_{i}$ denote the reward of the $i$-th item, and $cos(i,j)$ represent the cosine similarity between the $i$-th and $j$-th items in the selected subset. $W$ is a subset of $S$ that includes the most recently selected examples, reducing the emphasis on earlier selections. For example, if $w=10$, $W$ consists of the last 10 selected samples from $S$, while all previously selected examples are excluded from diversity considerations. The objective of MMR is to maximize rewards while ensuring sufficient diversity among the selected items. This iterative process continues until a termination criterion, such as reaching a predefined maximum number of tokens, is met.
$\operatorname*{argmax}_{i\in R}\left[\alpha\cdot r_{i}-(1-\alpha)\cdot\max_{j\in W}\text{cos}(i,j)\right]$ (1)
FPS
The concept of FPS originates from the field of 3D computer vision [12]. Its primary goal lies in selecting a diverse set of points from a given point cloud, which aids in hierarchical feature extraction for downstream applications. The process begins with a randomly selected initial point. In each subsequent iteration, a new point is chosen based on its distance from all previously selected points. When comparing FPS to MMR, we find that both are greedy methods that promote diversity by selecting points that differ from those chosen. However, FPS does not incorporate the concepts of a context window or reward. If we modify FPS to include these elements, the modified FPS will be equivalent to MMR, with the key difference being that MMR uses cosine similarity, while FPS relies on Euclidean distance for measuring similarity.
$\operatorname*{argmax}_{i\in R}\left[\max_{j\in S}\text{dist}(i,j)\right]$ (2)
2.2 Combine MMR and FPS with LLM for Diversity on Q&A and Summarization
Extending MMR and FPS techniques for LLMs in tasks such as Q&A and summarization is relatively straightforward as shown in Fig. 1. These techniques employ a greedy approach to iteratively balance the similarity of selected sentences or chunks to the query with the diversity among the selected sentences or chunks. This method enhances the likelihood of selecting the most relevant sentences or chunks for LLMs in downstream tasks. Lastly, inspired by [15], a heuristic rearrangement scheme is implemented to enhance the likelihood of identifying the correct answer from the retrieved documents.
Q&A
To evaluate the ability of LLMs on accurately extracting the correct answer, a query, a document, and a corresponding answer are initially provided. Documents are pre-processed by dividing them into sentences or chunks, and their embeddings are extracted beforehand. Both the query and the segmented documents are processed using encoder-only models to generate embeddings. In MMR, similarity is measured using the cosine angle, whereas in FPS, Euclidean distance is used to assess similarity. For benchmarking Q&A performance, two metrics should be evaluated:
  1. 1.
    Pre-LLM recall: whether the answer exists in the selected content before being sent to the LLM.
  2. 2.
    Post-LLM recall: whether the answer appears in the LLM’s output.
If the first metric shows significant improvement, the benefit of diversity becomes evident. Otherwise, the advantage of diversity may be limited. If the first metric improves while the second metric does not, it indicates that the performance of downstream tasks may be constrained by the capabilities of the LLM [15].
Summarization
In summarization tasks, datasets typically consist of a document paired with a corresponding golden summary created by experts. When no specific query is provided, the process begins by dividing the document into manageable chunks. Encoder-only models are employed to generate embeddings for these chunks, and the mean of these embeddings is used to represent the query embedding. Following this, the same methodology as in the previous Q&A task is applied to extract content that optimizes both reward and diversity.
The selected chunks are ordered to align with their original sequence in the document. These ordered chunks are sent to the LLM for summarization. We recognize that evaluating the extracted content before it is submitted to the LLM for summarization may not be particularly meaningful. Instead, we assess the quality of the LLM-generated summary by comparing it to the golden summary using metrics such as ROUGE [16] or LLM-as-a-judge [17].
Figure 2: The impact of different hyperparameters: $\alpha$, $w$, $c_{r}$ on the recall of the Natural Question dataset of single document Q&A. The first and second subfigures illustrate the recall ratios of answers contained in the selected documents for SB+MMR and SB+FPS. When the weight parameter $w=1$, they are equivalent to SB. From the results, we can conclude that both SB+MMR and SB+FPS outperform SB. The last two subfigures display the latency of SB+MMR and SB+FPS. SB+FPS shows slightly worse performances than SB+MMR, and the latency of SB+MMR is significantly lower, especially when the context window is very long. Considering these two aspects, SB+MMR is more suitable for practical use compared to SB+FPS.
Q&A Natural Question Trival Q&A Narrative Q&A
$c_{r}$=0.05 $c_{r}$=0.1 $c_{r}$=0.2 $c_{r}$=0.05 $c_{r}$=0.1 $c_{r}$=0.2 $c_{r}$=0.05 $c_{r}$=0.1 $c_{r}$=0.2
SB 46.28 58.60 69.41 63.44 71.64 78.33 18.60 21.04 25.61
SB+MMR 50.43 63.18 72.47 65.29 74.02 80.47 20.88 24.09 27.59
SB+FPS 50.88 63.23 72.33 65.25 73.18 80.07 21.34 24.24 27.44
Table 1: This table compares the performance of SB+MMR and SB+FPS against SB across three different datasets and three compression ratios, focusing on the recall of the correct answer within the selected documents.
Q&A Natural Question Trival Q&A Narrative Q&A
$c_{r}$=0.05 $c_{r}$=0.1 $c_{r}$=0.2 $c_{r}$=0.05 $c_{r}$=0.1 $c_{r}$=0.2 $c_{r}$=0.05 $c_{r}$=0.1 $c_{r}$=0.2
SB (index sort) 40.44 51.74 64.25 75.66 75.95 76.25 15.40 17.98 18.60
SB (sort) 40.25 50.81 60.61 75.76 76.15 76.38 13.72 15.55 17.07
SB (1:1) 40.25 51.55 60.24 75.29 76.33 76.47 14.18 16.31 17.99
SB (2:1) 41.28 50.99 61.08 75.31 76.33 76.87 14.18 16.46 17.84
SB (3:1) 39.41 51.09 60.05 75.56 76.13 76.57 14.63 16.01 17.23
SB+MMR (index sort) 45.76 57.25 67.90 75.73 76.20 76.72 18.60 19.05 20.27
SB+MMR (sort) 44.45 55.19 64.35 76.20 76.85 77.00 16.31 16.46 17.38
SB+MMR (1:1) 45.48 55.57 63.60 76.23 76.77 77.34 16.46 16.62 16.92
SB+MMR (2:1) 45.20 55.29 63.32 75.90 76.20 76.65 15.09 15.55 16.01
SB+MMR (3:1) 43.80 55.85 63.51 76.05 76.72 76.92 16.16 16.92 16.31
SB+FPS (index sort) 46.79 59.12 67.71 75.93 76.13 76.60 17.68 17.68 19.05
SB+FPS (sort) 45.76 57.25 63.32 75.83 76.45 76.87 16.31 16.62 17.23
SB+FPS (1:1) 46.14 57.90 62.76 75.78 76.35 77.09 16.62 17.07 16.77
SB+FPS (2:1) 45.76 56.13 62.20 75.88 76.23 77.02 15.85 16.62 16.46
SB+FPS (3:1) 44.27 55.10 63.69 76.40 76.77 76.95 15.70 16.46 16.01
Table 2: This table compares the performance of SB+MMR and SB+FPS against SB across three different datasets and three compression ratios, focusing on the recall of the correct answer within the LLM responses.
3 Experiments
The experiments conducted in this paper focus on three main topics: 1. Single Document Question Answering (Q&A), 2. Multiple Documents Question Answering (Q&A), and 3. Single Document Summarization.
For Single Document Q&A, the goal is to choose the correct answer from a set of candidate sentences within a single document. In Multiple Document Q&A, all documents in the dataset are firstly divided into chunks and then combined, and a query is used to find the correct answers across the entire dataset. Because the dataset size is too large, approximation methods are used to enhance efficiency and speed. Specifically, two metrics are evaluated: 1. recall of the correct answer in the extracted document, and 2. recall of the correct answer in the LLM response. The benefit of diversity is primarily reflected in the improvement of the first metric, while performance improvements in Q&A and summarization are mainly indicated by the second metric.
For summarization, various hyperparameters are considered in the optimization process:
  1. 1.
    The weight balance between reward and diversity, denoted as $\alpha$,
  2. 2.
    The context window size, $w$,
  3. 3.
    The compression ratio, $c_{r}$, or the maximum number of selected tokens, $T_{\text{max}}$.
3.1 Single Document Q&A
For single document Q&A, three datasets are included: 1. Natural Question [18], 2. Trival QA [19] and 3. Narrative QA [20]. For each dataset, it is composed of thousands of (query, document, answer) pairs where the answer exists within the document and answers the query. For each document, we split it into sentence using Spacy package [21]. SentenceBERT (SB) is utilized as the encoder to extract embeddings from sentences in different experiments [22]. Then, different compression ratios, i.e., $c_{r}=0.05,0.1,0.2$ are utilized. Here, $c_{r}=0.05$ represents that the fraction of the number of selected tokens over the total number of tokens should be 0.05, i.e., the termination condition when 0.05 of all tokens are selected for answering the query. As for $\alpha$ and $w$, different hyperparameters are tested in a two-level iteration. The first coarse-level iteration utilizes $\alpha$ from [0, 0.25, 0.5, 0.7, 0.8, 0.9, 0.95, 1] and $w$ from [0, 10, 100, 1000, 1000000]. Then, the best performing from the first coarse-level iteration is selected. For the second granular-level iteration, it further divides the neighbors of the best performing first coarse-level hyperparameters and select the ones that have the best performance. An example of the best performing hyperparameters in the Natural Question dataset in the coarse level can be found in Fig. 2. In particular, for Natural question, $\alpha=0.55$ and $w=1$ is the best for MMR and $\alpha=0.5$ and $w=1$ is the best for FPS. For Narrative Q&A, $\alpha=0.55$ and $w=5$ is the best for both MMR and FPS. For Trival Q&A, $\alpha=0.6$ and $w=3$ is the best for MMR and $\alpha=0.7$ and $w=1$ is the best for FPS.
Natural Question GPT4 GPT3.5
$T_{max}$=120k $T_{max}$=50k $T_{max}$=20k $T_{max}$=10k $T_{max}$=5k $T_{max}$=2k
E5 70.7 69.4 68.5 66.2 64.2 57.4
E5+MMR 71.5 71.5 69.8 67.2 65.1 57.8
Narrative Q&A GPT4 GPT3.5
$T_{max}$=120k $T_{max}$=50k $T_{max}$=20k $T_{max}$=10k $T_{max}$=5k $T_{max}$=2k
E5 13.42 10.06 6.7 4.88 4.88 4.57
E5+MMR 22.56 20.43 15.85 14.94 12.20 7.01
Trival Q&A GPT4 GPT3.5
$T_{max}$=120k $T_{max}$=50k $T_{max}$=20k $T_{max}$=10k $T_{max}$=5k $T_{max}$=2k
E5 84.62 81.15 74.01 70.24 65.08 56.55
E5+MMR 88.99 85.81 82.24 78.57 74.01 65.08
Table 3: This table compares the performance of E5+MMR against E5 across three different datasets, focusing on the recall of the correct answer within the selected documents.
Natural Question GPT4 GPT3.5
$T_{max}$=120k $T_{max}$=50k $T_{max}$=20k $T_{max}$=10k $T_{max}$=5k $T_{max}$=2k
E5 (index sort) 45.7 50.5 52.7 45.9 49.2 45.4
E5 (sort) 55.2 56.6 54.6 51.5 50.5 47.8
E5 (1:1) 53.5 55.9 55.8 50.5 50.9 46.9
E5 (2:1) 54.5 57.5 56.4 51 50.5 47.1
E5 (3:1) 54.7 57 54.8 51.3 49.8 47.4
E5+MMR (index sort) 46.2 50.3 53.2 47.6 50.9 48.6
E5+MMR (sort) 56.4 57.9 57 51.3 51.4 47.7
E5+MMR (1:1) 55 57.2 55.9 50.8 50.7 47.8
E5+MMR (2:1) 57.2 55.3 56.2 50.7 52.2 48.7
E5+MMR (3:1) 56.3 56.4 55.6 51.6 52.3 47.4
Table 4: This table compares the performance of E5+MMR against E5 on Natural Question, focusing on the recall of the correct answer within the LLM responses.
Narrative Q&A GPT4 GPT3.5
$T_{max}$=120k $T_{max}$=50k $T_{max}$=20k $T_{max}$=10k $T_{max}$=5k $T_{max}$=2k
E5 (index sort) 10.37 9.15 8.54 5.18 4.57 4.88
E5 (sort) 10.67 10.59 8.23 6.1 4.57 5.18
E5 (1:1) 9.76 9.45 7.93 4.88 4.88 5.18
E5 (2:1) 10.06 9.15 7.93 5.18 3.96 5.18
E5 (3:1) 10.98 9.45 7.93 5.79 4.57 4.88
E5+MMR (index sort) 10.67 10.37 11.28 4.88 6.1 4.88
E5+MMR (sort) 12.8 10.67 10.37 6.1 6.1 4.27
E5+MMR (1:1) 11.89 10.06 11.28 6.4 6.71 4.88
E5+MMR (2:1) 11.59 10.67 10.98 6.4 5.79 5.49
E5+MMR (3:1) 11.89 10.98 10.06 6.7 7.01 4.88
Table 5: This table compares the performance of E5+MMR against E5 on Narrative Q&A, focusing on the recall of the correct answer within the LLM responses.
Trival Q&A GPT4 GPT3.5
$T_{max}$=120k $T_{max}$=50k $T_{max}$=20k $T_{max}$=10k $T_{max}$=5k $T_{max}$=2k
E5 (index sort) 74.21 73.21 73.12 65.18 64.19 64.68
E5 (sort) 73.81 73.91 72.42 63.59 65.08 65.57
E5 (1:1) 74.4 73.12 72.92 64.29 64.29 64.98
E5 (2:1) 73.81 73.81 72.72 64.09 64.58 64.29
E5 (3:1) 73.31 74.11 72.72 63.99 64.38 64.78
E5+MMR (index sort) 74.7 74.9 73.51 66.47 66.87 64.88
E5+MMR (sort) 74.9 74.8 73.12 64.29 65.57 66.07
E5+MMR (1:1) 75.2 75.5 73.31 65.38 65.38 65.08
E5+MMR (2:1) 74.6 74.11 72.62 64.88 65.77 65.38
E5+MMR (3:1) 74.7 74.31 73.02 65.67 65.48 64.98
Table 6: This table compares the performance of E5+MMR against E5 on Trival Q&A, focusing on the recall of the correct answer within the LLM responses.
Based on the results across various datasets, we can assert that diversity significantly enhances the recall of the correct answer within the selected document, as demonstrated in Table 1, showing an improvement of 2% to 5%. When the extracted sentences are summarized by GPT4 using the prompt shown in Figure 3, the advantages of SB+MMR and SB+FPS over SB alone remain evident, as shown in Table 2. Additionally, we observe that the performance of Trivial Q&A after LLM is better than the retrieved sentences, with a consistent result of approximately 76%. This suggests that the performance is largely influenced by the LLM, possibly due to pretraining on Trivial Q&A, even when the retrieved documents are provided. FPS, using distance as the evaluation metric, performs slightly worse than MMR, which uses cosine similarity. Moreover, MMR is faster than FPS because computing cosine similarity is quicker than Euclidean distance in Python, especially as the compression ratio increases, as shown in Figure 2. This conclusion generally holds true across different datasets. The speed advantage of MMR becomes more critical as the number of candidates increases with the dataset size. Consequently, MMR will be used in the multiple document comparison in the next section.
Inspired by the paper "Lost in the Middle" [15], we sorted the selected sentences by different methods. The term "index sort" refers to sorting the sentences in their original order within the document. In comparison, "SB (m:n)" refers to allocating the first selected m sentences with highest scores at the beginning, the next n sentences with highest scores at the end, and then another m sentences at the beginning, continuing this pattern until all sentences are allocated. Specifically, "SB (sort)" is equivalent to "SB (1:0)" and does not alter the sequence of selected sentences. As shown in Table 2, SB (index sort) performs best because the original sequential information of the selected sentences in the document, despite missing some internal information, makes the most sense for GPT-4 in downstream tasks.
3.2 Mutiple Documents Q&A
For multiple documents Q&A, the same three datasets are utilized. In these datasets, the number of documents and the length of documents are relatively long, making it impractical to split each document into sentences. Instead, we follow the general framework of RAG to split each document into chunks of 512 tokens, with an overlapping ratio of 0.5 (i.e., 256 tokens) between any two adjacent chunks. To extract embeddings from these chunks and adhere to the standard pipeline of RAG, we apply the E5 model [23]. After applying the chunking strategy, the number of chunks can still reach nearly 1 million, which is impractical for exact search. To facilitate approximate search, principal component analysis (PCA) [24] is first applied to reduce the dimensionality of the embeddings, followed by clustering [25] to ensure the average number of chunks is less than 10k. Unlike single document Q&A, we set the maximum number of tokens rather than the compression ratio as the threshold for the maximum number of tokens selected. Specifically, $T_{max}$ is set to 2k, 5k, or 10k for GPT-3.5 and 20k, 50k, or 120k for GPT-4. Other settings remain the same. Different hyperparameters for $\alpha$ and $w$ are tested. For the Natural Questions dataset, $\alpha=0.9$ and $w=5$ yield the best results for GPT-3.5, while $\alpha=0.7$ and $w=5$ are optimal for GPT-4 in MMR. For Narrative Q&A, $\alpha=0.8$ and $w=30$ are best for GPT-3.5, and $\alpha=0.7$ and $w=300$ are best for GPT-4 in MMR. For Trivia Q&A, $\alpha=0.7$ and $w=20$ are best for GPT-3.5, and $\alpha=0.8$ and $w=300$ are best for GPT-4 in MMR. From the results, we observe that the optimal values for $\alpha$ and $w$ are generally larger for multiple document Q&A compared to single document Q&A.
When evaluating the performance of multiple-document Q&A systems, we observe a pattern similar to that of single-document Q&A. Specifically, the E5+MMR method shows a significant improvement over E5 in recall of the answers in retrieved documents, as demonstrated in Table 3, with a margin exceeding 10%. Additionally, E5+MMR outperforms E5 for post-LLM recall as shown in Tables 4, 5, and 6. However, future research should prioritize enhancing the LLM’s ability to utilize the retrieved documents effectively, rather than merely focusing on retrieving more accurate documents, as the LLM itself is the bottleneck. This observation is further corroborated in Trivial Q&A, where the results consistently achieve 64% accuracy for GPT3.5 and 76% for GPT4, irrespective of the retrieved document. Last, unlike single-document Q&A, placing important chunks at the beginning and ending positions of the prompt can provide benefits, particularly in Natural Question scenarios, as shown in Table 4, which can lead to a 10% improvement. This finding aligns with the conclusions of the paper "Lost in the Middle". The most relevant chunks to the query should be positioned either at the beginning or the end of the prompt.
SquAD Sentence Chunk size: 256 Chunk size: 512
10k 5k 2k 10k 5k 2k 10k 5k 2k
SB/E5 86.8 83.7 78.5 95 92.7 86.3 96.7 94.3 86.6
SB/E5+MMR 90.1 89.4 86.6 97 96.6 95.4 99 97.8 96.7
Table 7: This table compares the performances of sentence splitter and chunk splitter of size 256 and 512 on SquAD, focusing on the recall of the correct answer within the selected documents.
SquAD Sentence Chunk size: 256 Chunk size: 512
10k 5k 2k 10k 5k 2k 10k 5k 2k
SB/E5 (index sort) 70.4 73 71.7 78.2 81.1 79.2 80.2 82.9 78.2
SB/E5 (sort) 71.3 70.1 67.9 79.9 80 76.6 82.9 82.2 76.5
SB/E5 (1:1) 72 71.8 69.5 80.3 80.4 77.5 82.7 82 77.4
SB/E5 (2:1) 71.9 71.7 69.4 80.4 80.1 77.4 82.9 81.9 77.5
SB/E5 (3:1) 71.8 71 68.7 81.4 79.9 77.7 82.7 82.2 77.3
SB/E5+MMR (index sort) 68.7 74.2 75 76.5 81.4 83 76.1 84.2 84.9
SB/E5+MMR (sort) 71.5 71.8 72.5 82 81.9 82.6 84.1 84.5 84.7
SB/E5+MMR (1:1) 72.5 73.6 72.9 81.3 82.2 81.7 82.3 82.9 84.6
SB/E5+MMR (2:1) 71.9 73.1 73.2 81.8 82.5 83.1 84.5 84.5 85
SB/E5+MMR (3:1) 71.7 73.4 71.9 81.4 82.5 82.9 83.4 84.3 84.5
Table 8: This table compares the performances of sentence splitter and chunk splitter of size 256 and 512 on SquAD, focusing on the recall of the correct answer within the LLM responses.
Summarization Datasets gov_report legal
Rouge GPT4 WR Rouge GPT4 WR
SB 17.7 24.24 11.3 35.82
SB+MMR 18 $\frac{72.65+77.86}{2}=\textbf{75.26}$ 11.8 $\frac{71.64+56.72}{2}=\textbf{64.18}$
Table 9: This table compares the performance of SB+MMR against SB on gov_report and legal, using ROUGE and LLM-as-a-judge.
3.3 Sentence and Chunk Splitter Comparison on SquAD
For the comparison between sentence and chunk splitters on multiple documents Q&A, only the SQuAD dataset will be considered. The dataset sizes for Natural Questions, TriviaQA, and NarrativeQA are too large, making sentence-level experiments difficult. For the sentence-level splitter, Spacy is used. For the chunk-level splitter, a threshold of 256 or 512 tokens with 50% overlap between adjacent chunks is applied. All segmented sentences or chunks are mixed, reduced in dimension through PCA, and clustered for downstream tasks. Similar to previous experiments, $T_{max}$ is set to 2k, 5k, or 10k for GPT3.5. For the sentence-level splitter, the best parameters are $\alpha=0.25$ and $w=1000$. For the 256-token chunk-level splitter, the best parameters are $\alpha=0.5$ and $w=300$. For the 512-token chunk-level splitter, the best parameters are $\alpha=0.3$ and $w=300$. The results are consistent with previous findings. SB/E5+MMR significantly outperforms SB/E5, as shown in Table 7, with a 10% increase in recall of the correct answer within the selected documents. This recall increment of SB/E5+MMR over SB/E5 still exists in the LLM response, as shown in Table 8. "Index sort" generally performs better for sentence-level splitting, while sorting based on score is usually beneficial for chunking. A new takeaway is that chunk-level performance is better than sentence-level, with even better results for larger chunk sizes.
3.4 Single Document Summarization
For single document summarization, we include two datasets: the gov report [26] and legal documents [27]. We utilize GPT3.5 for summarization. To achieve this, we filter examples that are less than 15k tokens and then apply MMR to select sentences within each document until it reaches the predetermined threshold of 8k tokens. For the 1. gov report, the best parameters are $\alpha=0.9$ and $w=10$. For the legal documents, the best parameters are $\alpha=0.925$ and $w=300$. After selecting and ordering the selected sentences based on their original sequence, they are sent to GPT3.5 to generate the final summary using a specific prompt in Figure. 4. For both datasets, expert-written golden summaries are provided for each document. We evaluate the quality of generated summary using the ROUGE score by comparing with the golden summary. In addition, summaries by SB and SB+MMR are compared using LLM-as-a-Judge through GPT4. To address the position bias problem, we switch the sequences of the two summaries in two runs and average the win rate (WR). Our experiments reveal that diversity improves summary quality, as indicated by increased ROUGE scores and a higher LLM-as-a-Judge WR. Additionally, experiments on our internal data show that diversity is particularly beneficial for long emails, articles, and logs, where redundancy is a significant issue due to repetitive content, greetings, and long URLs. Diversity avoid overestimating information similar to the query.
4 Conclusion
This study proves the benefits of diversity through MMR and FPS to LLM performances on Q&A and summarization. From the retrival viewpoint, the recall is greatly improved both for sentence and chunk-level splitter, especially when $\alpha$ and $w$ are properly selected. This recall rate increment is maintained after LLM generation. However, future research should pay more attention to improve the LLM’s capability to find answers from the retrieved documents. MMR shows slightly better performances compared with FPS, and its latency property is much better, which greatly increase the potential of usage in application. For sentence-level splitter, arranging the selected sentences in their original sequence is usually beneficial and for chunk-level splitter, putting more important chunks at the beginning and ending positions are beneficial. Lastly, given a multiple document Q&A like SquAD, chunk-level splitter usually has a better performance compared with sentence-level splitter. Lastly, these conclusion on Q&A can be extended to summarization task.
5 Limitation
There are several limitation on this works. To begin with, we only work on English dataset, while multilingual datasets should be tested to prove the importance of diversity on other language. In addition, this work focuses on research dataset while more work is supposed to be conducted on industrial datasets. Lastly, for extremely large dataset, more engineering work on parallelization like tree structures should be conducted to reduce latency.
Appendix A Example Appendix
A.1 Prompts for Q&A and Summarization
Here are the prompts for Q&A in Figure. 3 and summarization in Figure. 4.
Figure 3: Prompts for Q&A
Figure 4: Prompts for Summarization